# LoRA Learns Less and Forgets Less

### v1

@@ C.1 MetaMathQA (Math IFT)
Both Self-Verification (Weng et al., 2022) and FOBAR (Jiang et al., 2024) fall under the category of “backward reasoning,” where the question starts with a given condition and requires reasoning backwards to solve for an unknown variable. In order to generate new mathematical questions, a numerical value in the original question is masked as a variable X, and the question is rephrased accordingly.
@@
  • FOBAR (55k samples, Jiang et al. (2024)): this approach is similar to Self-Verification but directly appends the answer to the question using the template “If we know the answer to the above question is $\hat{A}_{c}$, what is the value of unknown variable $x$?”
@@
Example C.1:  MetaMathQA type: GSM_Rephrased
@@
Example C.2:  MetaMathQA type: MATH_SV
@@
C.2 Magicoder-Evol-Instruct-110k (Code IFT)
Example C.3:  Magicoder-Evol-Instruct-110k instruction: Please amend the subsequent Python script so that it includes a ’while’ loop rather than the existing ’for’ loop, which iterates through the items of an integer list. The script currently has a bug where it attempts to print an object that is outside the bounds of the list. Fix this error and modify the script to use ’while’ instead of ’for’ loop. Ensure your script correctly handles empty lists.
@@ C.2 Magicoder-Evol-Instruct-110k (Code IFT)
C.3 Starcoder Python (Code CPT)
Example C.4:  Starcoder-Python
@@ C.3 Starcoder Python (Code CPT)
C.4 OpenWebMath (Math CPT)
Example C.5:  OpenWebMath url: http://math.stackexchange.com/questions/222974/probability-of-getting -2-aces-2-kings-and-1-queen-in-a-five-card-poker-hand-pa
@@ C.4 OpenWebMath (Math CPT)
Appendix D Theoretical Memory Efficiency Gains with LoRA for Single and Multi-GPU Settings
@@ D.1 Training on a Single GPU
D.2 Training on a Multiple with Fully Sharded Data Parallelism
Past approaches for training LLMs across multiple GPUs include model parallelism, where different layers of the LLM are stored on different GPUs. However this requires high communication overhead and has very poor throughput (Rajbhandari et al., 2020). Fully Sharded Data Parallelism (FSDP) shards the parameters, the gradient, and the optimizer states across GPUs. This incredibly efficient and actually is competitive with the memory savings offered by LoRA in certain settings.
FSDP sharding the parameter and optimizer states across N devices results in less memory usage relative to LoRA. LoRA on the other hand enables training on GPUs with far less memory and also emanes training without needing as many GPUs to shard across.
@@ D.2 Training on a Multiple with Fully Sharded Data Parallelism
Standard industry level GPUs have on-device memory between 16 GB (e.g. V100s) and 80 GB (e.g. A100s and H100s). As Table S1 demonstrates, the per-GPU memory requirements for training a 7B parameter model decrease drastically as the number of GPUs increase. The memory requirements for training a 7B model with Adam + LoRA on a single GPU are 15.12 GB, but the same per-GPU memory requirement for training a 7B model with Adam but without LoRA on 8 GPUs is 14 GB. In this 8 GPU scenario, the efficiency gains from LoRA disappear.
@@
We do the same analysis for a 405B parameter model to highlight how LoRA is beneficial as model size scales (Table S3).
@@
Table S1: Theoretical memory required to store the model and optimizer state during training for a 7B parameter model. Note that the numbers exclude memory needed to store activations. FSDP sharding the parameter and optimizer states across N devices results in less memory usage relative to LoRA. LoRA on the other hand enables training on GPUs with far less memory and also emanes training without needing as many GPUs to shard across.
@@
Table S3: Theoretical memory required to store the model and optimizer state during training for a 405B parameter model. Units are in GB

### v2

@@ G.1 MetaMathQA (Math IFT)
Both Self-Verification (Weng et al., 2022) and FOBAR (Jiang et al., 2024b) fall under the category of “backward reasoning,” where the question starts with a given condition and requires reasoning backwards to solve for an unknown variable. In order to generate new mathematical questions, a numerical value in the original question is masked as a variable X, and the question is rephrased accordingly.
@@
  • FOBAR (55k samples, Jiang et al. (2024b)): this approach is similar to Self-Verification but directly appends the answer to the question using the template “If we know the answer to the above question is $\hat{A}_{c}$, what is the value of unknown variable $x$?”
@@
Example G.1:  MetaMathQA type: GSM_Rephrased
@@
Example G.2:  MetaMathQA type: MATH_SV
@@
G.2 Magicoder-Evol-Instruct-110k (Code IFT)
As stated in the main text, this dataset contains 72.97M tokens of programming questions and answers (Wei et al., 2023). It reproduces the “Evol-Instruct” dataset of WizardCoder (Luo et al., 2023b) by iteratively prompting an LLM (GPT-4) to increase the difficulty of a set of question-answer pairs from Code Alpaca (Chaudhary, 2023). The dataset can be found here: https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K
Example G.3:  Magicoder-Evol-Instruct-110k instruction: Please amend the subsequent Python script so that it includes a ’while’ loop rather than the existing ’for’ loop, which iterates through the items of an integer list. The script currently has a bug where it attempts to print an object that is outside the bounds of the list. Fix this error and modify the script to use ’while’ instead of ’for’ loop. Ensure your script correctly handles empty lists.
@@ G.2 Magicoder-Evol-Instruct-110k (Code IFT)
G.3 Starcoder Python (Code CPT)
As stated in the main text, this dataset consists of permissively licensed repositories from GitHub, including Git commits, in 80+ programming languages (Li et al., 2023a) . We chose the Python subset and sub-sampled it to 20B tokens. The full dataset can be found here: https://huggingface.co/datasets/bigcode/starcoderdata
Example G.4:  Starcoder-Python
@@ G.3 Starcoder Python (Code CPT)
G.4 OpenWebMath (Math CPT)
As stated in the main text, this dataset contains 14.7B tokens derived from mathematical web pages from Common Crawl, correctly formatted to preserve mathematical content such as LaTeX equations (Paster et al., 2023) . The dataset can be found here: https://huggingface.co/datasets/open-web-math/open-web-math. As can be seen from the example below, this dataset contains a large amount of English.
Example G.5:  OpenWebMath url: http://math.stackexchange.com/questions/222974/probability-of-getting -2-aces-2-kings-and-1-queen-in-a- five-card-poker-hand-pa
@@ G.4 OpenWebMath (Math CPT)
There is often some confusion about the memory gains that vanilla LoRA offers both in theory and in practice. In Appendix H we discuss some of the theoretical benefits of LoRA, and show how it can enable training both on GPUs with less memory and on fewer total GPUs (in the multi-GPU setting). In Appendix I we show how LoRA in practice leads to memory savings relative to full finetuning, but can in fact lead to slower throughput for particular hardware and software settings.
Appendix H Theoretical Memory Efficiency Gains with LoRA for Single and Multi-GPU Settings
@@ H.1 Training on a Single GPU
H.2 Training on Multiple GPUs with Fully Sharded Data Parallelism
Past approaches for training LLMs across multiple GPUs include model parallelism, where different layers of the LLM are stored on different GPUs. However this requires high communication overhead and has very poor throughput (Rajbhandari et al., 2020). Fully Sharded Data Parallelism (FSDP) shards the parameters, the gradient, and the optimizer states across GPUs. This is incredibly efficient and is actually competitive with the memory savings offered by LoRA in certain settings.
FSDP sharding of the parameter and optimizer states across N devices results in less memory usage relative to LoRA. LoRA on the other hand enables training on GPUs with far less memory and also enables training on fewer GPUs.
@@ H.2 Training on Multiple GPUs with Fully Sharded Data Parallelism
Standard industry level GPUs have on-device memory between 16 GB (e.g. V100s) and 80 GB (e.g. A100s and H100s). As Table S14 demonstrates, the per-GPU memory requirements for training a 7B parameter model decrease drastically as the number of GPUs increases. The memory requirements for training a 7B model with Adam + LoRA on a single GPU are 15.12 GB, but the same per-GPU memory requirement for training a 7B model with Adam but without LoRA on 8 GPUs is 14 GB. In this 8 GPU scenario, the efficiency gains from LoRA disappear.
@@
We do the same analysis for a 405B parameter model to highlight how LoRA is beneficial as model size scales (Table S16). This is particularly relevant now that Llama-3-405B has been released by Meta (Dubey et al., 2024).
@@
Table S14: Theoretical memory required to store the model and optimizer state during training for a 7B parameter model. Note that the numbers exclude memory needed to store activations. FSDP sharding the parameter and optimizer states across N devices results in less memory usage relative to LoRA. LoRA on the other hand enables training on GPUs with far less memory and also enables training without needing as many GPUs to shard across.
@@
Table S16: Theoretical memory required to store the model and optimizer state during training for a 405B parameter model. Units are in gigabytes (GB)
Appendix I LoRA Throughput and Memory Measurements
We report training efficiency comparisons between full finetuning and models trained with LoRA for various choices of rank. We measured both the throughput (in tokens per second) and peak active memory (in GB) for training runs representative of the experiments reported in the paper. We performed the runs using a single node of 8×8$\times$8 ×H100-80GB GPUs. We used a per-GPU micro batch size of 1 and targeted all linear layer weights with LoRA (i.e. both Attention and MLP).
In Figure S10 we observe that there is a significant gap between full finetuning and LoRA runs, related to the additional overheads of the LoRA computations. In general, LoRA leads to an approximately $15\%$ reduction in throughput for a given batch size. LoRA with higher ranks is slower than lower ranks across all batch sizes; this is particularly noticeable for rank $r=512$. Similarly, LoRA settings with higher batch sizes have slightly higher throughput relative to lower batch sizes. Some of the slowdown is intrinsically related to the overheads of performing LoRA, since in practice it involves more computations of intermediate activations. However, we note that we did not optimize the LoRA implementation and used the publicly available HuggingFace peft library, which might be amenable to further optimizations that could reduce the gap in throughput.
For peak memory, we notice that for small batch sizes, LoRA provides a substantial reduction in peak memory ($\sim 40\%$). This is expected since the optimizer state is significantly smaller when using parameter efficient methods. However, as batch size increases, the size of intermediate activations increases proportionally, dominating the required memory. We limit the per GPU micro batch size to 8 to prevent out of memory errors, so for batch sizes 64 and above, we perform gradient accumulation. This leads to the throughput and memory stabilizing for batch size 64 and above, with just around ($\sim 15\%$ memory savings) for larger batch sizes.
Figure S10: Throughput and Memory Measurements for LoRA vs. full finetuning. (left) Training throughput measured in tokens per second across all 8 GPUs. (right) Peak active memory used by the training process in a single GPU (max GPU memory is 80GB).